# Superconducting spintronics with electron symmetry filtering and interfacial spin-orbit coupling

### v1

@@ Abstract
Over the recent years, crossroads of magnetism and superconductivity led to the emerging field of superconducting spintronics. A cornerstone of this venture is the generation of equal-spin triplet Cooper pairs in superconductor–ferromagnet hybrids, enabling long-range spin-polarized supercurrents and magnetic control over superconducting quantum states for the development of energy-efficient cryogenic devices. Until now, nearly all superconducting spintronic devices have relied on direct interfaces between superconductors and ferromagnets, since it was believed that an insulating barrier would decouple spin and charge transport. This assumption, however, appears to be invalid when a thin spin- and orbit-filtering barrier couples epitaxial ferromagnet and the superconductor. Symmetry filtering plays a crucial role in enhancing giant tunneling magnetoresistance (TMR) by selectively allowing specific electronic states to tunnel through the barrier. Such a mechanism is key for high-performance spintronic devices like magnetic random access memories, magnetic sensors or spin-light emitting diodes. On the other hand, spin-orbit coupling (SOC) is a central mechanism for perpendicular magnetic anisotropy in spintronics. Recently, it has become clear that SOC is crucial in mediating the interactions in heterostructures combining superconductors and ferromagnets, otherwise antagonistic materials where exotic interfacial quantum phenomena have been discovered over the last decade. Building on recent advances in studies of various V/MgO/Fe(100)-based, this manuscript provides a comprehensive review of superconducting spintronics driven by electron symmetry filtering and interfacial SOC. It emphasizes the critical role of a crystalline MgO barrier in selectively transmitting specific electronic states between V(100) and Fe(100). The manuscript also highlights how interfacial SOC enables symmetry mixing, allowing for the interaction between ferromagnetic and superconducting orderings though MgO(100). This mutual interaction, mediated by interfacial SOC, facilitates the conversion of spin-singlet to spin-triplet Cooper pairs. The work provides key insights into designing SOC based superconductor-ferromagnet hybrid structures for advanced superconducting spintronic functionalities.
@@ I Introduction
Although technically challenging, the growth and investigation of fully epitaxial S/F hybrids opens new avenues in superconducting spintronics, revealing physical mechanisms beyond those anticipated by existing theories. Most notably, epitaxy allows to introduce a new versatile tool: electron spin and orbital symmetry filtering. Symmetry-dependent spin-polarized currents enabled by epitaxial growth have been central to both the prediction and discovery of giant tunnelling magnetoresistance (TMR) in Fe/MgO/Fe tunnel junctions Butler2001; Mathon2001; Parkin2004; Yuasa2004. Incorporating symmetry filtering into superconducting spintronics can can significantly strengthen the influence of interfacial spin-orbit coupling (SOC) between the S and F order parameters. In the case of Fe/MgO interfaces, the symmetry filtering mechanism is known to yield exceptionally high spin polarization, up to $87\%$ Lu2009. Previously, symmetry filtering in S/F hybrids was only discussed in epitaxial Fe(100)/V(100) multilayer heterostructures in the context of the emergence of large MR near the superconducting critical temperature ($T_{C}$Moodera2008. Similar conclusions (although with smaller superconducting spin-valve effects) were reported later for V/Fe/V/Fe/CoO epitaxial multilayers Nowak2012.
@@
On top of this, perfect epitaxy has a number of crucial effects on the interaction between the F and S parts of hybrid devices: First, it may enhance electron interference effects in transport. Coherent phenomena such as MacMillan and Tomasch resonances have been used to demonstrate long-range penetration of LRTs into ferromagnets Visani2012; Costa2022. Second, it enables precise control over the magnetocrystalline anisotropy (MCA) of the F layer. This requires atomically sharp interfaces only achievable through epitaxial growth, which also facilitates the fine-tuning of interfacial spin textures, including perpendicular magnetic anisotropy Ikeda2010. This in turn strongly influences the coupling between the S and F order parameters. Third, epitaxial growth makes it feasible to insert thin insulating barriers (I) between the S and F layers, forming S/I/F structures. Though seemingly counterintuitive, these structures can host strong interfacial Rashba and Dresselhaus SOC. The reduced density of interfacial defects in these hybrids could help to distinguish the contributions of Rashba and Dresselhaus SOC Costa2021 and investigate shot nose. Last but not least, epitaxy offers superior homogeneity and reproducibility, critical both for meaningful comparisons with theoretical models (typically based on ideal, defect-free interfaces) and reliable fabrication of devices such as Josephson junctions and S/F/S-based qubits, where interfacial SOC is expected to play a pivotal role Cai2022.
@@
This manuscript presents detailed examination and discussion of current progress in normal and superconducting spintronics in epitaxial V/MgO/Fe based hybrids, with a focus on the crucial role of electron symmetry filtering and interfacial SOC in these systems. The current introductory Section I ends by motivating the work and with an exposition of the background that will be used during the rest of the review. Then, in Section II we provide a detailed explanation of the structure of the junctions under study and their fabrication process. Next, Section III examines the normal-state magnetic and transport properties, including ab-initio calculations of the interfacial SOC, the characterization of the junctions above $T_{C}$ and demonstrates how SOC already has a strong influence over the density of states (DOS) and transport. Section IV discusses techniques to characterize the SOC-induced coupling between the ferromagnetic and superconducting states. Specifically, we discuss the spin-triplet superconductivity-induced change in the MCA, the observation of magnetoanisotropic Andreev reflection (MAAR), and superconducting quasiparticle-induced electron-electron interference in the ferromagnetic electrodes. Section V summarizes some of the latest results in these type of junctions, which offer a distinct experimental evidence of LRTs, particularly through the observation of giant subgap shot noise. Finally, section VI offers an introduction into the current line of work: our first attempt to investigate electron transport in S/F/S junctions with lateral geometry, which could provide more direct evidence for the elusive LRTs by enabling the observation of a magnetization-controlled Josephson effect.
Background and motivation
@@
Spin-orbit coupling is an intra-atomic relativistic interaction between an electron’s spin and orbital motion, which plays a key role in solid-state systems Zutic2004. While the importance of SOC in the generation of spin currents in spintronics is now well established Jo2024, its role in generating spin supercurrents in S/F hybrid structures was only recognized in the past decade Bergeret2013; Banerjee2018; Martinez2020; Amundsen2024. In such S/F systems, SOC enables a range of phenomena not present in homogeneous materials, including the emergence of Majorana zero modes Manchon2015, the paramagnetic Meissner effect Espedal2016, strong anisotropy in low-bias conductance (Magneto Anisotropic Andreev Reflection MAAR) Hogl2015; Vezin2020; Martinez2020, and giant thermoelectric response feng-nanoenergy-24; ouassou-prb-22; GonzalezRuano2023; Tuero2025. It also influences the superconducting critical temperature, which becomes dependent on the magnetization orientation Banerjee2018, and could give rise to non-reciprocal transport effects such as the superconducting diode behavior spin-SDE; Ili2024. Moreover, SOC facilitates the generation of LRTs, significantly affecting spin dynamics in S/F hybrids Jeon2020. Spin orbit coupling has also been predicted to enable magnetization-controlled Josephson effects in S/F/S junctions, where it mediates the interaction between the superconductor and the ferromagnet Costa2017; eskiltprb2019; Bujnowski2019. Additionally, SOC could modify the MCA in epitaxial ferromagnets GonzalezRuano2020; GonzalezRuano2021. Since the SOC strength can be electrically tuned, it offers promising avenues for the advancement of superconducting spintronic devices and realizing topological quantum computing Cai2022.
@@
The main motivation behind the investigation of V/MgO/Fe-based epitaxial hybrid structures, as reviewed here, arises from three key features that distinguish these S/F hybrids from practically all previously studied superconducting spintronic systems. First, the Fe/MgO interface and, to a lesser extent, the V/MgO interface induce a strong SOC (primarily Rashba, due to the breaking of the translational symmetry in the atomic lattice at the interface). Second, the Fe/MgO system provides a spin polarization up to $80\%$, closely approaching the Fe electrode to a fully spin-polarized ferromagnet. Third, as discussed earlier, the symmetry mismatch between the electronic structures of V(100) and Fe(100) creates a scenario in which virtually all electrons transferred between the ferromagnet and superconductor experience SOC, a unique situation in superconducting spintronics. Electrons whose symmetry is transformed from $\Delta_{2}$ to $\Delta_{1}$ via SOC are efficiently transmitted through the MgO barrier. This symmetry selective transmission mechanism is similar to the well known underlying effect responsible for the record-high tunneling magnetoresistance (TMR) observed at room temperature in Fe/MgO/Fe magnetic tunnel junctions (MTJs)Yuasa2004; Parkin2004.
II Samples growth and characterization
@@
Four sets of epitaxial single-crystalline samples have been elaborated using Molecular Beam Epitaxy (MBE) in a chamber operating with a base pressure of approximately $1.1\texttimes\times 10^{-10}$ Torr. Sketches of their vertical structures are illustrated in Figure 1a-d, accompanied by the detailed multilayer structural sequence of each of the samples in Figure 1e.
@@
Figure 1: Sketch of the structure of the types of junctions studied. (a) F/F, (b) N(S)/F, (c) N(S)/F/F, (d) F/N(S)/F. The different colors represent each material (see legend). The Au top layer is not shown as it’s the same in all the samples. (e) Table with detailed information of the vertical structure of the samples.
@@
The details on the growth and characterization of the standard vertical MTJs (sample S0 which is Fe(40 nm)/MgO(12 ML)/Fe(10 nm)/Co(20 nm) could be found in Ref.Tiusan-2007.
@@
The first type of S/F hybrid samples (S1) consists of a V(40 nm)/MgO(12 ML)/Fe(10 nm) MTJ, with vanadium as the bottom electrode and iron as the top electrode (Figure 1b). The second samples (S2) are a double tunnel junctions comprising a bottom V electrode and two ferromagnetic Fe electrodes (middle and top), separated by single-crystal MgO barriers (Figure 1c). The third samples (S3) is a double-barrier MTJ with Fe as the bottom and top electrodes and V as the middle electrode (Figure 1d). In all three cases, the vanadium layers become superconducting at low temperatures ($\sim$ 4 K). All samples were grown on (100) oriented MgO substrates that were degassed at 650 °C for 30 minutes. During the subsequent growth of a 10 nm MgO seed layer, the substrate temperature was maintained at 450 °C. This seed layer serves to suppress residual carbon segregation at multilayer interfaces Tiusan-2007.
@@
For S1 and S2, a 2 nm Cr seed layer was deposited at $\sim$ 30 °C to improve the wetting and growth of the subsequent V(40 nm) electrode. This V layer was deposited at room temperature and subsequently annealed at 450 °C for 30 minutes to enhance crystallinity, as illustrated by the Reflection High-Energy Electron Diffraction (RHEED) images in Figure 2a. The MgO tunnel barrier was then epitaxially grown at $\sim$ 100 °C (Figure 2b), with its thickness precisely monitored using RHEED oscillations (Figure 2c). A 10 nm Fe electrode was deposited at the same temperature and annealed for crystallization and flattening at 400 °C for 20 minutes (Figure 2d). For S1, the growth was completed with a 15 nm Au capping layer.
@@
Figure 2: Monitoring and characterization of the growth procces of the sample S1. (a) Typical RHEED image for the annealed V (40nm) layer. (b) RHEED image. The yellow square marks the area averaged for the (c) intensity oscillatons during the growth of the MgO barrier on top of V are shown. (d) RHEED image for the Fe layer on top of the MgO barrier.
For S3 (Figure 1d), the multilayer growth started on an MgO(100)//MgO(10 nm) base template. A 45 nm Fe bottom electrode was deposited at $\sim$ 30 °C and annealed at 500 °C for 30 minutes to improve crystallinity and achieve an atomically flat surface. A first single crystal MgO barrier was then epitaxially grown layer by layer at $\sim$ 70 °C, with its thickness monitored via RHEED oscillations. On top of this barrier, a 40 nm V electrode was epitaxially grown at the same temperature and annealed for flattening at 450 °C for 30 minutes. A second single crystal MgO barrier was subsequently epitaxially deposited at $\sim$ 100 °C, followed by the growth of a 10 nm Fe electrode at $\sim$ 90 °C. This electrode was annealed at 400 °C for 20 minutes to promote crystallization and surface smoothing. Similarly to S2, the top Fe electrode was magnetically hardened by the deposition of a 20 nm Co overlayer at the same temperature. The final structure was capped with a 15 nm Au layer for protection.
Following the MBE growth, all samples were patterned into in square-shape junctions of a size from $10\times 10$ to $40\times 40$ $\mu{m^{2}}$ (with the diagonal along [100]) using UV lithography and ion etching, with the process monitored step by step by in-situ Auger spectroscopy (see Ref. Tiusan-2007 for more details).

### v2

@@ Abstract
Over the recent years, the crossroads of magnetism and superconductivity have led to the emerging field of superconducting spintronics. A cornerstone of this venture is the generation of equal-spin triplet Cooper pairs in superconductor–ferromagnet hybrids, enabling long-range spin-polarized supercurrents and magnetic control over superconducting quantum states for the development of energy-efficient cryogenic devices. Until now, nearly all superconducting spintronic devices have relied on direct interfaces between superconductors and ferromagnets, since it was believed that an insulating barrier would decouple spin and charge transport. This assumption, however, appears to be invalid when a thin spin- and orbit-filtering barrier couples an epitaxial ferromagnet and a superconductor. Symmetry filtering plays a crucial role in enhancing giant tunneling magnetoresistance (TMR) by selectively allowing specific electronic states to tunnel through the barrier. Such a mechanism is key for high-performance spintronic devices like magnetic random access memory, magnetic sensors or spin-light emitting diodes. On the other hand, spin-orbit coupling (SOC) is a central mechanism for perpendicular magnetic anisotropy in spintronics. Recently, it has become clear that SOC is crucial in mediating the interactions in heterostructures combining superconductors and ferromagnets, otherwise antagonistic materials where exotic interfacial quantum phenomena have been discovered over the last decade. Building on recent advances in studies of various V/MgO/Fe(100)-based systems, this manuscript provides a comprehensive review of superconducting spintronics driven by electron symmetry filtering and interfacial SOC. It emphasizes the critical role of a crystalline MgO barrier in selectively transmitting specific electronic states between V(100) and Fe(100). The manuscript also highlights how interfacial SOC enables symmetry mixing, allowing for the interaction between ferromagnetic and superconducting orderings through MgO(100). This mutual interaction, mediated by interfacial SOC, facilitates the conversion of spin-singlet to spin-triplet Cooper pairs. The work provides key insights into designing SOC-based superconductor-ferromagnet hybrid structures for advanced superconducting spintronic functionalities.
@@ I Introduction
Although technically challenging, the growth and investigation of fully epitaxial S/F hybrids opens new avenues in superconducting spintronics, revealing physical mechanisms beyond those anticipated by existing theories. Most notably, the epitaxy allows to introduce a new versatile tool: electron spin and orbital symmetry filtering. Symmetry-dependent spin-polarized currents enabled by epitaxial growth have been central to both the prediction and discovery of giant tunnelling magnetoresistance (TMR) in Fe/MgO/Fe tunnel junctions Butler2001; Mathon2001; Parkin2004; Yuasa2004. Incorporating symmetry filtering into superconducting spintronics can can significantly strengthen the influence of interfacial spin-orbit coupling (SOC) between the S and F order parameters. In the case of Fe/MgO interfaces, the symmetry filtering mechanism is known to yield exceptionally high spin polarization, up to $87\%$ Lu2009. Previously, symmetry filtering in S/F hybrids was only discussed in epitaxial Fe(100)/V(100) multilayer heterostructures in the context of the emergence of large magnetoresistance MR near the superconducting critical temperature ($T_{C}$Moodera2008. Similar conclusions (although with smaller superconducting spin-valve effects) were reported later for V/Fe/V/Fe/CoO epitaxial multilayers Nowak2012.
@@
On top of this, the perfect epitaxy has a number of crucial effects on the interaction between the F and S parts of hybrid devices. First, it may enhance electron interference effects in transport. Coherent phenomena such as MacMillan and Tomasch resonances have been used to demonstrate long-range penetration of LRTs into ferromagnets Visani2012; Costa2022. Second, it enables precise control over the magnetocrystalline anisotropy (MCA) of the F layer. This requires atomically sharp interfaces, only achievable through epitaxial growth, which also facilitates the fine-tuning of interfacial spin textures, including perpendicular magnetic anisotropy Ikeda2010. This in turn strongly influences the coupling between the S and F order parameters. Third, epitaxial growth makes it feasible to insert thin insulating barriers (I) between the S and F layers, forming S/I/F structures. Though seemingly counterintuitive, these structures can host strong interfacial Rashba and Dresselhaus SOC. The reduced density of interfacial defects in these hybrids could help to distinguish the contributions of Rashba and Dresselhaus SOC Costa2021 and investigate shot nose. Last but not least, the epitaxy offers superior homogeneity and reproducibility, critical for meaningful comparisons with theoretical models (typically based on ideal, defect-free interfaces) as well as for reliable fabrication of devices such as Josephson junctions and S/F/S-based qubits, where interfacial SOC is expected to play a pivotal role Cai2022.
@@
This manuscript presents a detailed examination and discussion of the current progress in normal and superconducting spintronics in epitaxial V/MgO/Fe-based hybrids, with a focus on the crucial role of electron symmetry filtering and interfacial SOC in these systems. The current introductory Section I ends by motivating the work and with an exposition of the background that will be used during the rest of the review. Then, in Section II we provide a detailed explanation of the structure of the junctions under study and their fabrication process. Next, Section III examines the normal-state magnetic and transport properties, including ab-initio calculations of the interfacial SOC, the characterization of the junctions above $T_{C}$ and demonstrates how SOC already has a strong influence over the density of states (DOS) and transport. Section IV discusses techniques to characterize the SOC-induced coupling between the ferromagnetic and superconducting states. Specifically, we discuss the spin-triplet superconductivity-induced change in the MCA, the observation of magnetoanisotropic low-bias conductance (MAAR), and superconducting quasiparticle-induced electron-electron interference in the ferromagnetic electrodes. Section V summarizes some of the latest results in these type of junctions, which offer a distinct experimental evidence of LRTs, particularly through the observation of giant subgap shot noise. Finally, Section VI offers an introduction into the current line of work: our first attempt to investigate electron transport in S/F/S junctions with lateral geometry, which could provide more direct evidence for the elusive LRTs by enabling the observation of a magnetization-controlled Josephson effect.
Superconducting spintronics with SOC
@@
Spin-orbit coupling is an intra-atomic relativistic interaction between an electron’s spin and its orbital motion, which plays a key role in solid-state systems Zutic2004. While the importance of SOC in the generation of spin currents in spintronics is now well established Jo2024, its role in generating spin supercurrents in S/F hybrid structures was only recognized in the past decade Bergeret2013; Banerjee2018; Martinez2020; Amundsen2024. In such S/F systems, SOC enables a range of phenomena not present in homogeneous materials, including the emergence of Majorana zero modes Manchon2015, the paramagnetic Meissner effect Espedal2016, strong anisotropy in low-bias conductance (magneto anisotropic Andreev reflection MAAR) Hogl2015; Vezin2020; Martinez2020, and giant thermoelectric response feng-nanoenergy-24; ouassou-prb-22; GonzalezRuano2023; Tuero2025. It also influences $T_{C}$, which becomes dependent on the magnetization orientation Buzdin1999; Banerjee2018, and could give rise to non-reciprocal transport effects such as the superconducting diode behavior spin-SDE; Ili2024. Moreover, SOC facilitates the generation of LRTs, significantly affecting spin dynamics in S/F hybrids Jeon2020. SOC has also been predicted to enable magnetization-controlled Josephson effects in S/F/S junctions, where it mediates the interaction between the superconductor and the ferromagnet Costa2017; eskiltprb2019; Bujnowski2019. Additionally, SOC could modify the MCA in epitaxial ferromagnets GonzalezRuano2020; GonzalezRuano2021. Since the SOC strength can be electrically tuned, it offers promising avenues for the advancement of superconducting spintronic devices and realizing topological quantum computing Cai2022.
@@
The main motivation behind the investigation of the V/MgO/Fe-based epitaxial hybrid structures, as reviewed here, arises from three key features that distinguish these S/F hybrids from practically all previously studied superconducting spintronic systems. First, the Fe/MgO interface and, to a lesser extent, the V/MgO interface induce a strong SOC (primarily Rashba, due to the breaking of the translational symmetry in the atomic lattice at the interface). Second, the Fe/MgO system provides a spin polarization up to $80\%$, closely approaching that of a fully spin-polarized ferromagnet in the Fe electrode. Third, as discussed earlier, the symmetry mismatch between the electronic structures of V(100) and Fe(100) creates a scenario in which virtually all electrons transferred between the ferromagnet and superconductor experience SOC, a unique situation in superconducting spintronics. Electrons whose symmetry is transformed from $\Delta_{2}$ to $\Delta_{1}$ via SOC are efficiently transmitted through the MgO barrier.
II Sample growth and characterization
@@
Four sets of epitaxial single-crystalline samples have been fabricated using molecular beam epitaxy (MBE) in a chamber operating with a base pressure of approximately $1.1\texttimes\times 10^{-10}$ Torr. Sketches of their vertical structures are illustrated in Figure 1a-d, accompanied by the detailed multilayer structural sequence of each of the samples in Figure 1e.
@@
Figure 1: Sketch of the structure of the types of junctions studied. (a) F/F, (b) N(S)/F, (c) N(S)/F/F, (d) F/N(S)/F. The different colors represent each material (see legend). The Au top layer is not shown as it is the same in all the samples. (e) Table with detailed information of the vertical structure of the samples.
@@
More details on the growth and characterization of the standard vertical MTJs (sample S0 which is Fe(40 nm)/MgO(12 monolayers)/Fe(10 nm)/Co(20 nm) could be found in Ref.Tiusan-2007.
@@
The first type of S/F hybrid samples (S1) consists of a V(40 nm)/MgO(12  monolayers)/Fe(10 nm) MTJ, with vanadium as the bottom electrode and iron as the top electrode (Figure 1b). The second samples (S2) are a double tunnel junctions comprising a bottom V electrode and two ferromagnetic Fe electrodes (middle and top), separated by single-crystal MgO barriers (Figure 1c). The third samples (S3) are double-barrier MTJs with Fe as the bottom and top electrodes and V as the middle electrode (Figure 1d). In all three cases, the vanadium layers become superconducting at low temperatures ($\sim$ 4 K). All samples were grown on (100)-oriented MgO substrates that were degassed at 650 °C for 30 minutes. During the subsequent growth of a 10 nm MgO seed layer, the substrate temperature was maintained at 450 °C. This seed layer serves to suppress residual carbon segregation at multilayer interfaces Tiusan-2007.
@@
For S1 and S2, a 2 nm Cr seed layer was deposited at $\sim$ 30 °C to improve the wetting and the growth of the subsequent V(40 nm) electrode. This V layer was deposited at room temperature and subsequently annealed at 450 °C for 30 minutes to enhance crystallinity, as illustrated by the reflection high-energy electron diffraction (RHEED) images in Figure 2a. The MgO tunnel barrier was then epitaxially grown at $\sim$ 100 °C (Figure 2b), with its thickness precisely monitored using RHEED oscillations (Figure 2c). A 10 nm Fe electrode was deposited at the same temperature and annealed for crystallization and flattening at 400 °C for 20 minutes (Figure 2d). For S1, the growth was completed with a 15 nm Au capping layer.
@@
Figure 2: Monitoring and characterization of the growth procces of the sample S1. (a) Typical reflection high-energy electron diffraction (RHEED) image for the annealed V (40nm) layer. (b) RHEED image. The yellow square marks the area averaged for the (c) intensity oscillatons during the growth of the MgO barrier on top of V are shown. (d) RHEED image for the Fe layer on top of the MgO barrier. All images are taken from [110] azimuth of the MgO substrate.
For S3 (Figure 1d), the multilayer growth started on an MgO(100)//MgO(10 nm) base template. A 45 nm Fe bottom electrode was deposited at $\sim$ 30 °C and annealed at 500 °C for 30 minutes to improve the crystallinity and achieve an atomically flat surface. A first single crystal MgO barrier was then epitaxially grown layer by layer at $\sim$ 70 °C, with its thickness monitored via RHEED oscillations. On top of this barrier, a 40 nm V electrode was epitaxially grown at the same temperature and annealed for flattening at 450 °C for 30 minutes. A second single crystal MgO barrier was subsequently epitaxially deposited at $\sim$ 100 °C, followed by the growth of a 10 nm Fe electrode at $\sim$ 90 °C. This electrode was annealed at 400 °C for 20 minutes to promote crystallization and surface smoothing. Similarly to S2, the top Fe electrode was magnetically hardened by the deposition of a 20 nm Co overlayer at the same temperature. The final structure was capped with a 15 nm Au layer for protection.
Following the MBE growth, all samples were patterned into in square-shape junctions of a size from $10\times 10$ to $40\times 40$ $\mu{\text{m}^{2}}$ (with the diagonal along the MgO substrate [100] direction) using UV lithography and ion etching, with the etching process being monitored step by step by in-situ Auger spectroscopy (see Ref. Tiusan-2007 for more details).